oatmeal5:

Part A: The area of a square is (4x2 − 12x + 9) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (16x2 − 9y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

Mercury:

part A: this is a perfect square trinomial in the form (a-b)^2 = a^2 - 2ab + b^2 comparing it to the expression (4x2 − 12x + 9), can you try to determine what your a and b terms would be? if b^2 = 9, what's b? if a^2 = 4x^2, what's a? same logic with part B, except this is a difference of squares (a-b)(a+b) = a^2 - b^2, again, identify your a and b terms